# Empirical distribution of ancestral lineages in populations with density-dependent interactions

### v2

@@ 4.2.1 Many-to-one formula in the large population limit
In this paragraph, we briefly introduce the homogeneous spine construction [4], as the proof of Theorem 4.5 builds on it. We first focus on populations of finite size. For $\psi=1$, individuals other than the spine behave exactly as in the original population process. Whenever the spine is of type $x$ in a population of type distribution $\mathbf{z}\in\mathcal{Z}$ such that $\mathbf{z}_{x}>0$, it reproduces to leave descendance $\mathbf{k}$ and become of type $y$ at rate $\mathbf{k}_{y}\tau_{\mathbf{k}}(x,\mathbf{z})$.
@@
$\left|\mathbb{E}_{\mathbf{z}^{K}}\left[\sum_{u\in\mathbb{G}(t)}F((x_{u}(s),Z^{K}(s))_{s\leq t})\right]-\sum_{x\in\mathcal{X}}\mathbf{z}_{x}\mathbb{E}_{x,\mathbf{z}}\left[\mathcal{W}(t)F((\Upsilon(s),z(s))_{s\leq t})\right]\right|\xrightarrow[K\to\infty]{}0.$ (22)
@@
$\mathbb{E}_{x,\mathbf{z}}[\mathcal{W}(t)F((\Upsilon(s),z(s))_{s\leq t})]=\sum_{x\in\mathcal{X}}\mathbf{z}_{x}\mathfrak{m}(x,\mathbf{z},t)\mathbb{E}_{x,\mathbf{z}}[F((\Upsilon^{(t)}(s),z(s))_{s\leq t})].$ (23)
@@
$\left|\mathbb{E}_{\mathbf{z}^{K}}\left[\sum_{u\in\mathbb{G}(t)}F((x_{u}(s),Z^{K}(s))_{s\leq t})\right]-\sum_{x\in\mathcal{X}}\mathbf{z}_{x}\mathfrak{m}(x,\mathbf{z},t)\mathbb{E}_{x,\mathbf{z}}[F((\Upsilon^{(t)}(s),z(s))_{s\leq t})]\right|\xrightarrow[K\to\infty]{}0.$ (24)
@@
$Y_{2}(t)$ $=\mathbf{e}(y_{0})+\sum_{(x,\mathbf{k},y)\in J^{*}}\int_{0}^{t}\int_{0}^{+\infty}\mathbf{1}_{\left\{Y_{2}(s-)=x,\theta\leq\mathbf{k}_{y}\tau_{\mathbf{k}}(x,\zeta_{1}(s-))\right\}}(\mathbf{e}(y)-\mathbf{e}(x))Q_{x,\mathbf{k},y}(ds,d\theta).$
@@
$\mathbb{P}(T_{K}\leq t_{K})\leq$ $\;\mathbb{P}(A_{K},\sum_{(x,\mathbf{k},y)\in J^{*}}\int_{0}^{t}\int_{0}^{+\infty}\mathbf{1}_{\left\{\mathcal{E}_{x,\mathbf{k},y}(\theta,Y_{1}(s-),\zeta_{1}(s-))\Delta\mathcal{E}_{x,\mathbf{k},y}(\theta,Y_{1}(s-),\zeta_{2}(s-))\right\}}Q_{x,\mathbf{k},y}(ds,d\theta)\geq 1)$ (26)
@@
where the last inequality follows from Lemma 4.7. Recall that $\lambda$ is bounded. Thus $t\mapsto\exp\left(\int_{0}^{t}\lambda(\Upsilon_{1}(s),z_{1}(s))ds\right)$ is Lipschitz continuous on $[0,T]$, for a Lipschitz constant which can be chosen to be independent from $x$ and $\mathbf{z}$. Hence
@@ 4.2.2 Speed of convergence
Let $t>0$ and $F\in\mathfrak{L}$. There exists $C>0$ such that for any sequence $(\varepsilon_{K})_{K\geq 0}$ of positive real numbers, for any $K\geq 1$, letting $\delta_{K}=C(\varepsilon_{K}+K^{-1}\varepsilon_{K}^{-1})$,
@@
Combining Equation (22) with Proposition 4.6 and 4.12 for $\varepsilon_{K}=K^{-1/2}$ finally yields Theorem 4.5.
@@
$\mathbb{P}\left(\sup_{s\in[0,t]}\|\zeta^{K}(s)-z(s)\|_{1}\geq\varepsilon_{K}\right)\leq\frac{C(t)}{K\varepsilon_{K}}.$
@@
As the family $(M^{K}_{x,\mathbf{k}})_{x,\mathbf{k}}$ is independent, $M^{K}$ thus is itself a square integrable martingale, and Doob’s inequality shows that
$\mathbb{E}\left[\sup_{s\in[0,t]}M^{K}(s)\right]\leq 4\mathbb{E}[\langle M^{K}\rangle(t)]=4\sum_{(x,\mathbf{k})\in J}\mathbb{E}[\langle M^{K}_{x,\mathbf{k}}\rangle(t)].$ (30)
Since $J$ is a finite set, it follows that there exists a non-negative, finite constant $c=\sum_{x,\mathbf{k}}c(x,\mathbf{k})$ such that, for any $K\geq 1$,
$\mathbb{E}\left[\sup_{s\in[0,t]}M^{K}(s)\right]\leq\frac{tc}{K}.$
@@
Combining Equations (30), (31) and (32) yields the existence of finite, non-negative constants $c_{1}$ and $c_{2}$ such that, for any $K\geq 1$,
$\mathbb{E}\left[\sup_{s\in[0,t]}\|\zeta^{K}(s)-z(s)\|_{1}\right]\leq\frac{tc_{1}}{K}+c_{2}\int_{0}^{t}\mathbb{E}\left[\sup_{u\in[0,s]}\|\zeta^{K}(u)-z(u)\|_{1}\right]ds.$
@@
$\mathbb{E}\left[\sup_{s\in[0,t]}\|\zeta^{K}(s)-z(s)\|_{1}\right]\leq\frac{C(t)}{K}.$
@@
Let $t>0$. Under Assumption 4.1, there exists a constant $C(t)>0$ such that for every $K\geq 1$ and $\varepsilon_{K}>0$, letting $\alpha_{K}=(K\varepsilon_{K})^{-1}$,
@@ 5 Discussion
Second, under appropriate assumptions, we have focused on sampling in the deterministic large population limit, and quantified the associated approximation error. In particular, the process describing the descendance of the spine then corresponds to a time-inhomogeneous multi-type branching process whose reproduction rates depend on a changing environment, given by $z$. If we assume that $z$ admits a stable equilibrium, then starting from this equilibrium, the descendance is described by a classical multi-type branching process [12].
@@
with $\varepsilon_{K}$ and $\alpha_{K}$ converging to zero and $t_{K}$ growing to infinity, as $K$ grows large. We consider expecting such control to be reasonable. Indeed, it corresponds to understanding and controlling the fluctuations of the finite-population process $\zeta^{K}$ around its deterministic limit, which is a well studied question with several classical regimes: Gaussian fluctuations for $\varepsilon_{K}=O(K^{1/2})$, which are related to the diffusion approximation, moderate deviations with $\varepsilon_{K}=O(K^{p})$ for $p\in(0,1/2)$ and large deviations where $\varepsilon_{K}=O(1)$ [11, 25, 27]. In particular, moderate and large deviations appear to be interesting regimes, as they allow to consider longer time scales. Nevertheless, in the case of the spinal population process, they are not immediate, due to a boundary problem arising from the fact that the spinal individual becomes negligible in large populations. As a consequence, these considerations are left for a futur work.
@@ Acknowledgments.
Let us turn to the case $\{T_{1}\leq h<T_{2}\}$. For $\mathbf{k}\in\mathbb{N}^{d}$ and $y\in\mathcal{X}$, we can compute the probability that the spine leaves descendance $\mathbf{k}$ and becomes of type $y$ at $T_{1}$, and does not reproduce anymore until time $s+h$:

### v3

@@ 4.2.1 Many-to-one formula in the large population limit
In this paragraph, we briefly introduce the homogeneous spine construction [4], as the proof of Theorem 4.5 builds on it. We first focus on populations of finite size. For $\psi=1$, individuals other than the spine behave exactly as in the original population process. Whenever the spine is of type $x$ in a population of type distribution $\mathbf{z}\in\mathcal{Z}$ such that $\mathbf{z}_{x}>0$, it is replaced by offspring $\mathbf{k}$ and becomes of type $y$ at rate $\mathbf{k}_{y}\tau_{\mathbf{k}}(x,\mathbf{z})$.
@@
$\left|\frac{1}{K}\mathbb{E}_{\mathbf{z}^{K}}\left[\sum_{u\in\mathbb{G}(t)}F((x_{u}(s),Z^{K}(s))_{s\leq t})\right]-\sum_{x\in\mathcal{X}}\mathbf{z}_{x}\mathbb{E}_{x,\mathbf{z}}\left[\mathcal{W}(t)F((\Upsilon(s),z(s))_{s\leq t})\right]\right|\xrightarrow[K\to\infty]{}0.$ (22)
@@
$\mathbb{E}_{x,\mathbf{z}}[\mathcal{W}(t)F((\Upsilon(s),z(s))_{s\leq t})]=\mathfrak{m}(x,\mathbf{z},t)\mathbb{E}_{x,\mathbf{z}}[F((\Upsilon^{(t)}(s),z(s))_{s\leq t})].$ (23)
@@
$\left|\frac{1}{K}\mathbb{E}_{\mathbf{z}^{K}}\left[\sum_{u\in\mathbb{G}(t)}F((x_{u}(s),Z^{K}(s))_{s\leq t})\right]-\sum_{x\in\mathcal{X}}\mathbf{z}_{x}\mathfrak{m}(x,\mathbf{z},t)\mathbb{E}_{x,\mathbf{z}}[F((\Upsilon^{(t)}(s),z(s))_{s\leq t})]\right|\xrightarrow[K\to\infty]{}0.$ (24)
@@
$Y_{2}(t)$ $=\mathbf{e}(y_{0})+\sum_{(x,\mathbf{k},y)\in J^{*}}\int_{0}^{t}\int_{0}^{+\infty}\mathbf{1}_{\left\{Y_{2}(s-)=x,\theta\leq\mathbf{k}_{y}\tau_{\mathbf{k}}(x,\zeta_{2}(s-))\right\}}(\mathbf{e}(y)-\mathbf{e}(x))Q_{x,\mathbf{k},y}(ds,d\theta).$
@@
$\mathbb{P}(T_{K}\leq t)\leq$ $\;\mathbb{P}(A_{K},\sum_{(x,\mathbf{k},y)\in J^{*}}\int_{0}^{t}\int_{0}^{+\infty}\mathbf{1}_{\left\{\mathcal{E}_{x,\mathbf{k},y}(\theta,Y_{1}(s-),\zeta_{1}(s-))\Delta\mathcal{E}_{x,\mathbf{k},y}(\theta,Y_{1}(s-),\zeta_{2}(s-))\right\}}Q_{x,\mathbf{k},y}(ds,d\theta)\geq 1)$ (26)
@@
where the last inequality follows from Lemma 4.7. Recall that $\lambda$ is bounded. Thus $t\mapsto\exp\left(\int_{0}^{t}\lambda(\Upsilon_{1}(s),z_{1}(s))ds\right)$ is Lipschitz continuous on $[0,T]$, for a Lipschitz constant which can be chosen to be independent of $x$ and $\mathbf{z}$. Hence
@@ 4.2.2 Speed of convergence
Let $t>0$ and $F\in\mathfrak{L}$. There exists $C>0$ such that for any sequence $(\varepsilon_{K})_{K\geq 0}$ of positive real numbers, for any $K\geq 1$, letting $\delta_{K}=C(\varepsilon_{K}+K^{-1/2}\varepsilon_{K}^{-1})$,
@@
Combining Equation (22) with Proposition 4.6 and 4.12 for $\varepsilon_{K}=K^{-1/4}$ finally yields Theorem 4.5.
@@
$\mathbb{P}\left(\sup_{s\in[0,t]}\|\zeta^{K}(s)-z(s)\|_{1}\geq\varepsilon_{K}\right)\leq\frac{C(t)}{\sqrt{K}\varepsilon_{K}}.$
@@
$M^{K}$ thus is itself a square integrable martingale, and Doob’s inequality shows that there exists a positive constant $\kappa$ such that
$\mathbb{E}\left[\sup_{s\in[0,t]}(M^{K}(s))^{2}\right]\leq\kappa\sum_{(x,\mathbf{k})\in J}\mathbb{E}[\langle M^{K}_{x,\mathbf{k}}\rangle(t)].$ (30)
Since $J$ is a finite set, it follows that there exists a positive, finite constant $c=\kappa\sum_{x,\mathbf{k}}c(x,\mathbf{k})$ such that, for any $K\geq 1$,
$\mathbb{E}\left[\sup_{s\in[0,t]}(M^{K}(s))^{2}\right]\leq\frac{tc}{K}.$
Noticing that $M^{K}\geq 0$ almost surely, Cauchy-Schwarz inequality finally leads to
$\mathbb{E}\left[\sup_{s\in[0,t]}M^{K}(s)\right]\leq\sqrt{\frac{tc}{K}}.$
@@
Combining Equations (30), (31) and (32) yields the existence of finite, non-negative constants $c_{1}(t)$ and $c_{2}$ such that, for any $K\geq 1$,
$\mathbb{E}\left[\sup_{s\in[0,t]}\|\zeta^{K}(s)-z(s)\|_{1}\right]\leq\frac{c_{1}(t)}{\sqrt{K}}+c_{2}\int_{0}^{t}\mathbb{E}\left[\sup_{u\in[0,s]}\|\zeta^{K}(u)-z(u)\|_{1}\right]ds.$
@@
$\mathbb{E}\left[\sup_{s\in[0,t]}\|\zeta^{K}(s)-z(s)\|_{1}\right]\leq\frac{C(t)}{\sqrt{K}}.$
@@
Let $t>0$. Under Assumption 4.1, there exists a constant $C(t)>0$ such that for every $K\geq 1$ and $\varepsilon_{K}>0$, letting $\alpha_{K}=(\sqrt{K}\varepsilon_{K})^{-1}$,
@@ 5 Discussion
Second, under appropriate assumptions, we have focused on sampling in the deterministic large population limit, and quantified the associated approximation error. In particular, the process describing the spine’s offspring then corresponds to a time-inhomogeneous multi-type branching process whose reproduction rates depend on a changing environment, given by $z$. If we assume that $z$ admits a stable equilibrium, then starting from this equilibrium, the offspring is described by a classical multi-type branching process [12].
@@
with $\varepsilon_{K}$ and $\alpha_{K}$ converging to zero and $t_{K}$ growing to infinity, as $K$ grows large. We consider expecting such control to be reasonable. Indeed, it corresponds to understanding and controlling the fluctuations of the finite-population process $\zeta^{K}$ around its deterministic limit, which is a well studied question with several classical regimes: Gaussian fluctuations for $\varepsilon_{K}=O(K^{-1/2})$, which are related to the diffusion approximation, moderate deviations with $\varepsilon_{K}=O(K^{-p})$ for $p\in(0,1/2)$ and large deviations where $\varepsilon_{K}=O(1)$ [11, 25, 27]. In particular, moderate and large deviations appear to be interesting regimes, as they allow to consider longer time scales. Nevertheless, in the case of the spinal population process, they are not immediate, due to a boundary problem arising from the fact that the spinal individual becomes negligible in large populations. As a consequence, these considerations are left for a futur work.
@@ Acknowledgments.
Let us turn to the case $\{T_{1}\leq h<T_{2}\}$. For $\mathbf{k}\in\mathbb{N}^{d}$ and $y\in\mathcal{X}$, we can compute the probability that the spine is replaced by offspring $\mathbf{k}$ and becomes of type $y$ at $T_{1}$, and does not reproduce anymore until time $s+h$: